find the roots of the quadratic equation 2x^2-5x+3=0
Answer by ewatrrr(24785)

 
Hi,
find the roots of the quadratic equation
2x^2-5x+3=0 = (2x-3)(x-1) = 0 x = 3/2 or x = 1
checking our factoring using FOIL
F First terms 2x^2
O Outside terms -2x
I Inside terms -3x
L Last terms 3